# Accelerating Mobile Language Model via Speculative Decoding and NPU-Coordinated Execution

### v3

@@ 2.3. Challenges: Aligning Retrieval-based Speculative Decoding with NPU
$\bullet$ NPU execution: static compute graph design hinders performance improvement for both prefill and decoding. Static compute graph constraints fixed input and output tensor shapes during execution. To support variable-length sequences in LLM inference, existing systems adopt a chunking strategy, i.e., splitting inputs into fixed-size segments (Xu et al., 2025b; Chen et al., 2025a). However, a unified chunk size for both prefill and decoding yields suboptimal performance. Figure 5 compares prefill and speculative decoding speed using different compute graphs on a Redmi K70 Pro smartphone, with a fixed acceptance ratio of 4.2 (based on evaluation results, see Section 4.5). Results show that compute graphs with larger chunk size improve prefill throughput but hurt decoding performance. For instance, despite a fast prefill speed (¡8 ms/token) of Qwen2.5 1.5B using graph with chunk size 32, the decoding speed using the same graph drops to about 160 ms/token. This tradeoff necessitates specialized compute graphs to improve performance for both prefill and decoding. However, on resource-constrained devices, naively maintaining multiple graphs in memory is costly, highlighting the needs of lightweight graph scheduling algorithms.
@@
Figure 6. Comparison of verification speed with different draft lengths, and the length distribution of drafts generated by existing R-SD method (Saxena, 2023) (histogram).
$\bullet$ Draft construction: divergence between task context and LLM output reduces the efficiency of speculative decoding. As discussed in Section 2.1, contexts and predictions often exhibit lexical gaps (i.e., normalized Levenshtein distance >0.11>0.11). These gaps result in frequent rejection of retrieved drafts, despite their success in capturing the similar semantics as LLM. This degrades decoding to a token-by-token process where NPUs offer little speedup due to underutilization. Lexical gaps arise from divergence between the task context and the LLM’s distribution, necessitating alignment of the two distributions.
$\bullet$ Draft verification: existing R-SD methods underutilize NPUs’ weight-stationary design due to drafting short sequences. The design of mobile NPUs, while ideal for compute-intensive operations such as prefill, is less effective when input sequences become short in the decoding stage. Existing R-SD methods perform retrieval by strict suffix matching between context and output (Somasundaram et al., 2024; He et al., 2024; Hu et al., 2024), which frequently results in short or empty drafts. Figure 6 demonstrates the influence of draft length on verification speed and the length distribution of generated draft on a summarization dataset (Xia et al., 2024) . Results show that over 90% of drafts fall below the NPU-unfriendly length, significantly reducing the verification speed as NPU computing capacity is underutilized. This results in a similar speed to CPU (shown in the horizontal line). For instance, the verification speed of Qwen2.5 1.5B drops from 5.7 ms/token to 51.1 ms/token when the draft length decreases from 32 to 1.
@@ 3.1. Overview
Figure 7. Overview of sd.npu framework.:
Design goal. sd.npu accelerates context-augmented generation by NPU-optimized retrieval-based speculative decoding. This entails (1) deploying and scheduling two compute graphs optimized for prefill and decoding, (2) aligning task context with LLM distribution, and (3) enriching drafts with sufficient high-value tokens to fully utilize NPU resources.
Workflow. Figure 7 illustrates the system architecture and workflow of sd.npu.
  • Offline stage. sd.npu partitions the LLM into multiple blocks, each containing several transformer layers. It profiles the load and compute latency of each block through warm-up to determine scheduling strategy.
  • Online stage. Upon receiving a request, sd.npu fetches task-relevant database and performs prefill using a SOTA chunk-based approach (Xu et al., 2025b; Chen et al., 2025a), splitting the input into chunks processed by a long-sequence optimized graph. During prefill, progressive graph scheduling (Section 3.2) incrementally switches the prefill-optimized graph to the decoding-optimized graph with load overlapped by computation. After prefill, output logits are used for in-context distribution calibration (Section 3.3), augmenting the database with calibrated contexts. sd.npu then performs retrieval-based speculative decoding using the decoding-optimized graph and the calibrated database. It iteratively constructs a drafting tree and verifies it via a confidence-based strategy to identify and reuse valuable tokens (Section 3.4). After an end-of-sequence token is predicted, the final response is returned and stored in the database manager as a historical context for future requests.
Key techniques. sd.npu incorporates three techniques to align retrieval-based speculative decoding with NPU offloading:
  • Progressive NPU graph scheduling (Section 3.2) enables asynchronous graph switching by overlapping load and computation across chunked prefill rounds.
  • In-context distribution calibration (Section 3.3) aligns task context with model expectations using prefill logits.
  • NPU-optimized draft reuse (Section 3.4) maximizes NPU utilization and reduces verification costs by reusing plausible tokens from rejected drafts.
@@ 3.2. Progressive Graph Scheduling
Challenges of fixed-shape compute graphs. As discussed in Section 2.3, mobile NPUs require fixed-shape compute graphs, forcing the prefill input to be chunked for variable-length requests. This constraint also forces prefill and decoding to share the same graph despite their distinct input patterns: prefill processes long sequences (e.g., chunk size 256), while decoding handles short inputs (e.g., chunk size 32). Using a unified graph thus leads to inefficiency. To address this, we consider two specialized graphs: $G^{1}$ optimized for prefill and $G^{2}$ optimized for decoding. We denote the computation latency of a forward pass of model $\operatorname{P}$ on graph $G^{x}$ ($x\in\{1,2\}$) as $T_{\text{compute}}(\operatorname{P};G^{x})$.
@@
Switching graph progressively. sd.npu gradually switches $G^{1}$ to $G^{2}$ during prefill. The design builds on two insights:
  • Graph equivalence by repetition: executing $G^{2}$ multiple times on shorter inputs yields the same result as one execution of $G^{1}$ on longer inputs, allowing early switching. As sd.npu maximizes draft lengths (Section 3.4), $G^{2}$ will not be too short (e.g., chunk/2 or chunk/4 of $G^{1}$), which will not incur much performance degradation.
  • Chunk-level scheduling: multiple block computation can be overlapped with a single block loading.
@@
Figure 8. Progressive graph scheduling example with $N=6$ and two chunks. (a) shows the workflow of chunk 1 computation; (b) compares prefill pipelines under different scheduling schemes.
Figure 8 illustrates progressive scheduling with $N=6$ and two input chunks. We suppose that sd.npu decides to overlap loading a block by two blocks of computation based on the profiling results. Initially, all blocks are loaded with $G^{1}$. sd.npu computes $\operatorname{P}_{1},\operatorname{P}_{2}$ with $G^{1}$ while loading $G^{2}$ for $\operatorname{P}_{6}$. Then $\operatorname{P}_{3},\operatorname{P}_{4}$ execute while $G^{2}$ loads for $\operatorname{P}_{5}$. To compute $\operatorname{P}_{5},\operatorname{P}_{6}$ under $G^{2}$, chunk 1 is split into sub-chunks matching $G^{2}$’s shape, with $G^{2}$ executed repeatedly. Prefill of chunk 1 finishes with output logits while half of the model is switched. The same procedure applies to chunk 2. In contrast, synchronous loading waits for all compute to finish, and naive asynchronous loading only overlaps in chunk 2, both incurring higher latency (Figure 8(b)).
Problem Formulation. Formally, given $N$ model blocks that sequentially execute prefill for $C$ chunks, sd.npu aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round. The objective is to minimize the overall latency, which is equal to the maximum of the prefill completion time and the switch completion time.
Scheduling methodology. This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (Korf, 1998) (proofed in Appendix A). sd.npu addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute. This strategy enables fully masking the load overhead by computation to minimize switching overhead. sd.npu finds the smallest $k$ by estimating $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})$ and $\widehat{T}_{\text{compute}}(\operatorname{P}_{i};G^{x}_{i})$ based on the profiling results. To hide load latency, the sum of the next $k$ blocks’ compute time must be larger than the load overhead of the last block not switched, formulated as:
@@ 3.3. In-Context Distribution Calibration
Challenges of language divergence. As discussed in Section 2.1, leveraging similarity on mobile LLMs is hindered by distribution divergence between task context and LLM modeling. Models often interpret the context with different lexical patterns, thus reducing the acceptance ratio of drafts directly generated from the context. These lexical gaps appear prominently at the token level. For example, in Figure 9, a user request states “… Clinton’s security detail has arrived at Des Moines …,” while the LLM, tasked with summarizing, prefers “Clinton arrived at Des Moines.” For classical retrieve-based speculative decoding approaches (Saxena, 2023; He et al., 2024; Hu et al., 2024), draft “Clinton’s security” is rejected due to lexical mismatch between token ’s and target arrived. This results in only one token is generated (arrived) by one forward pass as the following draft security is discarded, reducing decoding efficiency. Such mismatch problem arises in various mobile scenarios such as auto-reply or RAG-based QA, where human-written contexts diverge from the model’s distribution.
Figure 9. Overview of in-context distribution calibration with a summarization example. :
To tackle this issue, one may forcibly accept drafts with lexical gaps based on semantic-level similarity. However, setting a stringent similarity threshold results in a low acceptance ratio with only marginal performance gains, whereas a loose threshold causes non-negligible accuracy loss, e.g., drops from 0.9 to 0.75 if candidates are forcibly accepted ignoring the distribution divergence (Holsman et al., 2025; Bachmann et al., 2025). This necessitates calibration of draft and model distributions. A straightforward solution is to precompute a calibrated database by running large volumes of user inputs through the LLM, which suffers from poor generalization to unseen queries and requires prohibitive storage and offline processing. Another method is to fine-tune the LLM to adapt to specific task scenarios, which also consumes lots of resources and energy and is hard to cover the wide diversity of mobile tasks.
Leveraging prefill logits. To avoid heavy offline or fine-tuning costs, our key observation is that after prefill completes, next-token logits for each token in the entire context have already been calculated. These logits reflect how LLM encodes task semantics and can be used to approximate the model distribution. To that end, we can build a calibrated database for drafting with negligible performance overhead. Therefore, we propose in-context distribution calibration, an online, lightweight method that dynamically aligns context with the LLM distribution using prefill logits.
Workflow. After prefill, sd.npu first collects predicted logits for all positions. Then sd.npu builds a calibrated token tree representing the model distribution by depth-first search, which includes two steps:
@@
Overhead analysis. The main costs come from sampling and storing calibrated results. In our experiments, sampling incurs an amortized latency of about 2 ms per input token, while building and maintaining calibrated tokens peak below 500 MB memory for a 3B model (Section 4.3).
3.4. NPU-Optimized Draft Reuse
Challenges of short potentially acceptable drafts. Existing studies on R-SD focus on improving draft quality, proposing techniques such as pruning draft tree by estimating token confidence (Somasundaram et al., 2024; Hu et al., 2024; Wu et al., 2025a). However, as discussed in Section 2.3, a large proportion of drafts generated by existing methods have short lengths, which is harmful to the NPU efficiency. Naively lengthening drafts by loosing retrieving conditions cannot obtain performance benefits as irrelevant tokens will be rejected in verification.
Reuse plausible drafts. As highlighted in Section 2.1, the model output usually shares similar semantics with the task context despite existing lexical gaps. Although these gaps cause rejection due to token-level misalignment, the rejected tokens include correct semantics and may be accepted in the next verification. sd.npu identifies and reuses these plausible tokens to construct NPU-friendly drafts lengthened by task-relevant information.
@@ 3.4. NPU-Optimized Draft Reuse
Rejection causes analysis. We first analyze the rejection reasons to explore how to identify plausible tokens. Let $\mathbf{x}=[T,x_{1},\dots,x_{n}]$ be the draft sequence where $T$ is the most recently accepted token, and $\mathbf{y}=[y_{1},\dots,y_{n+1}]$ be the model’s predictions. Suppose the expected sequence is $\mathbf{z}=[z_{1},\dots,z_{n+1}]$ and rejection occurs at position $e$ due to a mismatch: $x_{e}\neq y_{e}$. Thus, the accepted sequence is $x_{1},\dots,x_{e-1},y_{e}$. According to the patterns of the expected sequence $z_{e}\dots z_{n}$, we categorize reuse opportunities into three general cases and report their distribution, which is obtained by manually labeling the rejection causes from a user case of summarization. Table 2 summarizes the results, showing that nearly 40% of the cases include reused sequence portions.
  • Missing: the expected sequence includes additional tokens $z_{e}\dots z_{e+\delta}\ (\delta\geq 0)$ not present in the draft, while the remaining draft tokens $x_{e}\dots x_{e+\epsilon}\ (\epsilon\geq 1)$ match the expectation. Thus, the expected sequence is $z_{e}\dots z_{e+\delta}x_{e}\dots x_{e+\epsilon}$. If the combined sequence $z_{e+\delta}x_{e}\dots x_{e+\epsilon}$ does not exist in the database, next retrieval will fail with no drafts generated. This could be recovered by reserving $x_{e}\dots x_{e+\epsilon}$.
  • Synonym: the draft includes a semantically equivalent but lexically different segment $x_{e}\dots x_{e+\gamma-1}\ (\gamma\geq 1)$, replaced in the expected response by $z_{e}\dots z_{e+\delta}$. Similar failure occurs if the post-synonym tokens $x_{e+\gamma}\dots x_{e+\epsilon}$ cannot be retrieved from database using $z_{e+\delta}$. In this case, $x_{e+\gamma}\dots x_{e+\epsilon}$ should be preserved.
  • Redundant: the entire draft $x_{e}\dots x_{n}$ is incorrect and fully diverges from the expected path. This is the only case where discarding the sequence is appropriate.
Reuse methodology. To leverage reusable tokens in the first two cases, sd.npu aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
@@
(3) $\gamma,\epsilon=\underset{\gamma,\epsilon}{\text{argmax}}(\epsilon-\gamma)\quad\text{s.t. }y_{i}=x_{i},\ \forall\ i\in[e+\gamma,e+\epsilon]$
Each reused segment is reserved for at least one iteration, and is discarded when the cumulative draft length of current decoding step exceeds a predefined threshold.
By maximizing draft length with potentially acceptable tokens, this reuse mechanism directly enhances NPU utilization and reduces verification times. The overhead of sd.npu’s drafting and reusing processes is near ignorable with ¡3 ms per decoding step, as discussed in Section 4.4.

### v4

@@ 3.1. Overhead of Static Graph Management
Figure 4. Latency breakdown of RAG decoding phase.
Consequently, runtimes are compelled to perform graph switching, i.e., unloading the prefill graph and reloading a decoding-specialized graph. However, this operation is far from free. Our breakdown of decoding latency (Figure 4) reveals that graph switching alone accounts for 29.0% of the total time for RAG tasks with 1K context. Naive switching blocks execution and saturates memory bandwidth, while keeping both graphs resident is infeasible due to the tight RAM budget of mobile devices. In contrast, CPU decoding avoids this overhead but suffers from significantly lower raw compute capability.
[…]
3.2. Ineffectiveness of Speculative Decoding on Mobile NPUs
@@
Figure 5. Comparison of verification speed with different draft lengths, and the length distribution of drafts generated by existing SD method (work:pld) (histogram).
Figure 6. Workflow of speculative decoding and an example of existing inefficiencies: a large proportion of computation is wasted or underutilized.
Besides graph scheduling challenge, the NPU‑only decoding phase faces another bottleneck: memory bandwidth and NPU underutilization. Generating a single token on the NPU is memory-bound due to low arithmetic intensity. In this regime, the NPU’s massive matrix array is starved, degrading throughput to CPU levels while consuming the fixed overhead of NPU kernel invocation. To amortize these overheads and saturate the NPU, the workload must be transformed from sequential vector operations into parallel matrix operations.
Speculative decoding (work:sd_is_lossless) is a widely adopted method to effect this transformation. SD uses a lightweight model (i.e., model-based methods (work:medusa; work:eagle)) or database (i.e., retrieval-based methods (work:pld; work:rest)) to generate multiple draft tokens and verifies them in parallel. In verification, only correct tokens are accepted and any tokens after the first mismatch is discarded (Figure 6a). This draft-then-verify schema allows SD to convert memory-bound decoding into compute-bound verification while maintaining lossless generation quality (work:sd_is_lossless). Among different approaches, model-based SD is infeasible for mobile devices as it requires a separate draft model (often >1>1GB), which competes for scarce RAM and compute capacity with the main LLM and the OS. Instead, for RAG workloads, the retrieved documents intrinsically contain the information needed for generation. Using the context as a draft source (work:sam) incurs negligible overhead, making it the ideal candidate for mobile RAG. Consequently, this paper adopts retrieval-based speculative decoding as the foundational mechanism for mobile acceleration.
Breakdown of SD inefficiency. Despite its theoretical suitability, we find that off-the-shelf algorithms fail to accelerate NPU inference. To understand this inefficiency, we decompose NPU computation during verification into three parts: (i) valid, resources spent on accepted drafts; (ii) wasted, resources spent on rejected drafts; and (iii) underutilized, the NPU’s idle capacity when the workload is too small. As show in Figure 6b, our measurements reveal that wasted (22%) and underutilized portions (75%) dominate the use of compute resources, leaving the effective throughput surprisingly low.
Figure 7. Effective throughput comparison. For NPU, throughput is dictated by accept length.
Root cause: constant-cost trap. This inefficiency stems from the unique performance characteristics of mobile NPUs. Unlike GPUs where cost scales linearly, NPU verification latency remains nearly constant up to a hardware saturation point $M$ (Figure 7). In this regime, decoding throughput is dictated solely by the total accepted length per step. To achieve speedup, the system must maximize the number of accepted tokens within this constant-time window. However, existing retrieval-based methods fail to exploit this characteristic. For instance, as shown in § 5.5, PLD (work:pld) achieves an average accepted length of only 2.53 tokens per step, resulting in less speedup. This failure manifests in two specific ways:
Challenge#2.1: wasted computation due to low draft acceptance. To preserve lossless generation, SD rejects drafts that lexically mismatch the base model’s predictions, even if they are semantically aligned. In RAG workloads, this lexical divergence is common, causing frequent rejections. Since the NPU incurs a near-constant high cost for each verification step regardless of the outcome, every rejection squanders a full pass of NPU computation. This directly contributes to a high percentage of wasted computation.
Challenge#2.2: underutilized computation due to short workloads. The NPU’s constant-cost model means that peak efficiency is only achieved with long drafts that saturate its compute units. However, existing SD methods for RAG often produce very short drafts. Empirically, over 90% of drafts contain fewer than 32 tokens (Figure 5). Verifying such short sequences leaves most of the NPU’s matrix units idle, paying the high cost of a full verification step for only a few tokens. This severely throttles effective throughput, sometimes making NPU even slower than CPU.
Together, these observations motivate the need for an new SD algorithm explicitly co-designed with RAG workloads and NPU execution characteristics.
@@ 4.1. Overview
Figure 8. Overview of sd.npu.
Design goal. To enable efficient RAG inference with end-to-end NPU offloading, sd.npu is designed as a runtime system that couples speculative decoding with NPU execution. Its key idea is a system–algorithm co-design: at the system level, sd.npu reduces graph-switching overhead, and at the algorithmic level, it introduces an NPU-optimized speculative decoding scheme.
Architecture and workflow. Figure 8 illustrates the overall architecture. It intercepts the standard prefill-decoding pipeline and orchestrates speculative decoding workloads tailored for NPU execution with the following workflow:
$\bullet$ Offline initialization. sd.npu partitions the LLM into several blocks, each containing a contiguous set of Transformer layers. A one‑time warm‑up profiling characterizes each block’s latency and memory footprint under different chunk sizes. This profile is used to determine the optimal partitioning and graph combination for the device’s NPU configuration. sd.npu also maintains a token database that includes chat records and relevant corpus to facilitate draft generation.
$\bullet$ Online execution. When an augmented user request arrives, sd.npu performs the following stages:
Prefill. The request is partitioned into fixed‑size chunks following the approach of prior mobile LLM systems (work:mllm; work:heterollm). NPU executes these chunks using a long‑sequence optimized (prefill) graph. Progressive graph scheduling4.2). During prefill, the system gradually transitions to the decoding‑optimized graph through block-by-block switching, overlapping I/O costs with prefill computation (§4.2). Distribution calibration. After prefill, the system collects the output logits and applies in‑context distribution calibration (§4.3) to augment the draft database. This procedure aims to reduce rejection-caused recomputation by aligning drafts’ token statistics with the model’s conditional distribution. Speculative decoding. With the draft database, sd.npu performs speculative decoding using the decoding‑optimized graph. It iteratively constructs tree-structured drafts and verifies them via a confidence-based policy that identifies and reuses valuable tokens (§4.4). When an end‑of‑sequence token is verified, sd.npu returns the merged output and commits the interaction into the database to facilitate future generation.
@@ 4.2. Progressive Graph Scheduling
Challenges of fixed-shape compute graphs. As discussed in §3.1, mobile NPUs require static compute graphs, forcing prefill and decoding to share the same graph despite their distinct input patterns: prefill processes long sequences (e.g., input length 256), while decoding handles short inputs (e.g., input length 1). Using a shared graph thus leads to inefficiency, while keeping two graphs resident in memory is wasting limited RAM budget. To address this, we consider switching between two specialized graphs: $G^{1}$ optimized for prefill and $G^{2}$ optimized for decoding. We denote the computation latency of a forward pass of model $\operatorname{P}$ on graph $G^{x}$ ($x\in\{1,2\}$) as $T_{\text{compute}}(\operatorname{P};G^{x})$.
@@
Progressive switching. Instead, sd.npu amortizes the heavy switching cost across the entire prefill through two insights:
$\bullet$ Graph equivalence: executing $G^{2}$ multiple times on shorter inputs is mathematically equivalent to a single execution of $G^{1}$ on longer inputs, allowing early switching. As sd.npu maximizes draft lengths (§4.4), $G^{2}$ will not be too short (e.g., chunk/4 of $G^{1}$), avoiding performance degradation.
$\bullet$ Chunk decomposition: through fine-grained block partitioning and scheduling, the I/O payload of a single block can be masked by several chunked prefill iterations.
@@
Figure 9. Progressive graph scheduling example with $N=6$ and two chunks. (a) workflow of computing chunk 1; (b) pipelines comparison under different scheduling.
Figure 9 illustrates progressive scheduling with $N=6$ and two input chunks. We suppose that sd.npu decides to overlap loading a block by two blocks of computation based on the profiling results. Initially, all blocks are loaded with $G^{1}$. sd.npu computes $\operatorname{P}_{1},\operatorname{P}_{2}$ with $G^{1}$ while loading $G^{2}$ for $\operatorname{P}_{6}$. Then $\operatorname{P}_{3},\operatorname{P}_{4}$ execute while $G^{2}$ loads for $\operatorname{P}_{5}$. To compute $\operatorname{P}_{5},\operatorname{P}_{6}$ under $G^{2}$, chunk 1 is split into sub-chunks matching $G^{2}$’s shape, with $G^{2}$ executed repeatedly. Prefill of chunk 1 finishes with output logits while half of the model is switched. The same procedure applies to chunk 2. In contrast, synchronous loading waits for all compute to finish, and naive asynchronous loading only overlaps in chunk 2, both incurring higher latency, as shown in Figure 9(b).
Problem Formulation. Formally, given $N$ model blocks (decided by offline profiling) that sequentially execute prefill for $C$ chunks, sd.npu aims to determine the time to load each block, i.e., choosing blocks $\{\operatorname{P_{i},\dots}\}$ to overlap load overhead of $\operatorname{P_{j}}$ by their computations at each compute-load round. The objective is to minimize the overall latency, which is equal to the maximum of the prefill completion time and the switch completion time.
Scheduling methodology. This scheduling problem is NP-hard as it is essentially equivalent to the 2-Partition problem (work:partition). sd.npu addresses this via a greedy algorithm, choosing the least $k$ consecutive blocks $\{\operatorname{P}_{i},\dots,\operatorname{P}_{i+k-1}\}$ that overlap a block load to execute. This strategy enables fully masking the load overhead by computation to minimize switching overhead. sd.npu finds the smallest $k$ by estimating $\widehat{T}_{\text{load}}(\operatorname{P}_{j};G^{2})$ and $\widehat{T}_{\text{compute}}(\operatorname{P}_{i};G^{x}_{i})$ based on the profiling results. To hide load latency, the sum of the next $k$ blocks’ compute time must be larger than the load overhead of the last block not switched, formulated as:
@@ 4.3. In-Context Distribution Calibration
Wasted computation due to rejection. To achieve lossless acceleration, speculative decoding rejects all draft tokens after the first incorrect token. Each rejection forces recomputation and falling back to token-by-token decoding, neutralizing speculative parallelism and wasting NPU cycles. This verification strategy requires lexical identity, resulting that semantically valid drafts are rejected. This lexical mismatches frequently arise in mobile RAG workloads such as summarization and QA, where retrieved text originates from human-written content with phrasing styles that deviate from the model’s learned patterns. As shown in Figure 10, given a draft “Clinton’s security detail has arrived at Des Moines,” the model’s own distribution predicts “Clinton arrived at Des Moines.” Although the semantics align, lexical mismatch at the token level (e.g., ’s vs. arrived) triggers rejection (work:pld; work:rest; work:sam). The verification then produces only one accepted token (arrived), forcing recomputation for the rest (security detail). Across NPU pipelines, this leads to redundant graph executions and degraded decoding throughput.
Figure 10. Overview of in-context distribution calibration with a summarization example.
Design goal. We aim to reduce wasted NPU computation caused by draft rejection, without compromising semantic accuracy. The key idea is to align drafts with the model’s intrinsic distribution such that they better match what the model is likely to verify. A straightforward solution is to precompute a calibrated database by running large volumes of user inputs through the LLM, which suffers from poor generalization to unseen queries and requires prohibitive storage and offline processing. Another method is to fine-tune the LLM to adapt to specific scenarios, which also consumes lots of resources and is hard to cover the wide diversity of RAG tasks.
Leverage prefill logits. During prefill, NPUs already compute next-token logits for all context tokens as part of the model’s forward pass. These logits encode the LLM’s internal probability distribution and naturally reveal how the model interprets the current task context. By reusing these prefill logits, we can estimate the model’s local token transition space and calibrate the retrieval distribution without additional forward passes or fine-tuning. We call this mechanism in-context distribution calibration, an online, lightweight technique that exploits prefill logits to dynamically align drafts with model-side token probabilities, effectively avoiding wasteful computation on rejected drafts.
Workflow. After prefill, sd.npu first collects predicted logits for all positions. Then sd.npu builds a token-level calibration structure that approximates the model’s local generation distribution by 2-step depth-first search:
@@
4.4. NPU-Optimized Draft Extension
Short drafts underutilize NPU. NPUs are optimized for dense, long-running workloads that amortize weight loading and maximize matrix reuse across many tokens. However, existing workloads are typically short: most drafts contain 2-8 tokens, well below the length needed to saturate systolic arrays, as shown in Figure 5. Simply lengthening drafts is ineffective because irrelevant tokens are immediately rejected, forcing recomputation and wasting NPU cycles. Therefore, extending drafts with plausible tokens is crucial for efficient NPU execution.
Observation: rejected tokens are not always incorrect. While SD rejects drafts with lexical mismatches, these tokens often encode semantically correct continuations. For example, when summarizing “Clinton’s security detail has arrived at Des Moines,” the model may reject “security” due to token-level mismatch, yet the phrase “Clinton arrived at” appears again in subsequent verification. Thus, the earlier rejected segment partially overlaps with future correct output, representing “wasted” computation that could be reused to extend future drafts and reduce underutilized NPU cycles.
@@ 4.4. NPU-Optimized Draft Extension
Empirical analysis of reuse opportunities. We analyze rejection causes by comparing drafts $\mathbf{x}=[T,x_{1},\dots,x_{n}]$ with model predictions $\mathbf{y}=[y_{1},\dots,y_{n+1}]$ and expected sequences $\mathbf{z}=[z_{1},\dots,z_{n+1}]$. Suppose a rejection occurs at position $e$ where $x_{e}\neq y_{e}$. The accepted prefix is $x_{1},\dots,x_{e-1},y_{e}$. Our manual annotation across a summarization dataset (Table 2) reveals that nearly 40% of rejected tokens appear again within subsequent decoding steps, indicating that they can be safely reused to extend NPU workloads. Specifically, we categorizes three major rejection patterns in Table 2:
$\bullet$ Missing: The expected output includes additional tokens absent from the draft, but subsequent draft tokens remain valid. Reusing $x_{e}\dots x_{e+\epsilon}$ preserves the valid suffix.
$\bullet$ Synonym: The rejected tokens are semantically equivalent but lexically different (e.g., “car” vs. “vehicle”). Retaining the following segment $x_{e+\gamma}\dots x_{e+\epsilon}$ allows reusing once the synonym mismatch resolves.
$\bullet$ Redundant: The entire segment diverges from the target sequence and must be discarded.
Reuse strategy. sd.npu dynamically identifies reusable token segments from previously rejected drafts and appends them to new drafts to form NPU-friendly workloads. Formally, sd.npu aims to: (1) determine the start and end positions $\gamma,\epsilon$ of the maximum reusable segment; (2) estimate the reuse lifetime $\delta$, i.e., how many iterations the reused segment should be reserved for verification.
@@
(3) $\gamma,\epsilon=\underset{\gamma,\epsilon}{\text{argmax}}(\epsilon-\gamma)\quad\text{s.t. }y_{i}=x_{i},\ \forall\ i\in[e+\gamma,e+\epsilon].$
This reused segment is retained across at least one decoding step and discarded once its accumulated lifetime or total draft length exceeds a predefined threshold. Thus, sd.npu effectively reuse wasted tokens and reduce underutilized ones, saving 70.5% of token computation without sacrificing accuracy. The additional overhead is discussed in § 5.4.